The length of a room is 5 less than twice the width. The perimeter is 80 feet.
Answer by anantha(86)

sol:
let width be x
twice the width=2x
length of the room=5 less than twice the width=2x-5
perimeter of the rectangle=2*(length+width)=2*(2x-5+x)=2*(3x-5)
according to the problem perimeter of the rectangle=80 feet
2*(3x-5)=80
dividing both sides by 2
2*(3x-5)/2=80/2
3x-5=40
adding both sides by 5
3x-5+5=40+5
3x=45
dividing by 3 on both sides
3x/3=45/3
x=15
width=15 feet
and length=2x-5=(2*15)-5=30-5=25 feet
check this problem:
perimeter =2*(length+width)=2(25+15)=2*40=80 feet
which is given